OpenStudy (josee):

I have no clue on how to do this.. it says let f(x)= 3x. find the number delta > 0 so that 0<|x-2|< delta => |f(x)-6| < epsilon... then it has a) epsilon=0.5 b)epsilon=0.1 c) epsilon=0.01 and d) epsilon is a (presumably small) positive number... :/

OpenStudy (anonymous):

So epsilon is given in each problem, so you need to come up with a formula for delta in terms of epsilon. So lets look at the expression: \[|f(x)-6|<\epsilon \]what we want to do is make the left hand side look like |x-2|. So doing a little bit of algebra we get: \[|f(x)-6|<\epsilon \iff |3x-6|<\epsilon \iff 3|x-2|<\epsilon \]\[|x-2|<\frac{\epsilon}{3}\] So let delta be epsilon over 3.

OpenStudy (josee):

Ok ok.. so whats the next step? I kinda dont understand what im lookin for

OpenStudy (anonymous):

So now we have a formula for delta in terms of epsilon. It is: \[\delta =\frac{\epsilon}{3}\] So for example, in a), the problem is, "find the number delta if epsilon is .5" The correct delta would be: \[\delta = \frac{.5}{3}=.16666\]

OpenStudy (josee):

OMG thank you so much!